# Infrared Small Target Detection in Satellite Videos: A New Dataset and A Novel Recurrent Feature Refinement Framework

### v2

@@ III-B2 Target Appearance Simulation
Due to remote sensing imaging system and optical diffraction effect, infrared small targets always appear as small points or diffraction spots. As shown in Fig. 3 (a), targets can be approximated by Gaussian kernels $G(h,w,\sigma)$ [70, 71, 59], where $h$, $w$ are major & minor axis, and $\sigma$ represents sigma value. We randomly simulate $N$$\in$$[1,6]$ Gaussian targets for a image sequence, and $h$$\in$$[1,9]$, $w$$\in$$[1,9]$, $\sigma$$\in$$[0.1,1]$. To generate the target appearance sequence of the $n^{th}$ target $G_{nt}(h_{t},w_{t},\sigma_{t})$, as shown in Fig. 1 (b), we randomly set $K$$\in$$[2,5]$ Gaussian kernels $G_{nt_{1}}$, $G_{nt_{2}}$,…,$G_{nt_{K}}$ ($t_{1}$=$1$, $t_{K}$=$T$, $\{t_{2}$$t_{K-1}\}$$\in$$(1,T)$) as the key target appearances in a sequence, and employ quadratic interpolation1 to generate intermediate appearances.
@@ III-B3 Target Trajectory Simulation
Figure 4 shows some examples of target trajectories in satellite video sequences [72]. It can be observed that real target trajectory is continuous and relatively smooth over a short time interval, which can be approximated using low-order curve with order $p$$\in$$(0,3)$ [59]. In addition, target may swerve to change the moving direction and velocity. To simulate the $n^{th}$ target trajectory with $S$$\in$$[0,2]$ swerving times, as shown in Fig. 1 (c), we first synthetic $S$ local target trajectories by low-order curves, and sequentially connect them end-to-end. Then, we employ 1D interpolation to smooth each intersection of two adjacent trajectories for a global smooth trajectory $(x_{nt}^{T},y_{nt}^{T})$. Figure 5 shows the trajectories of two swerved targets in the $1^{st}$ frame reference. It can be observed that target trajectories are locally continuous and globally smooth, which are in accordance with the real-world target motion. It is worth noticing in Fig. 5 (b1) that, the moving direction of one target is opposite to that of background, which indicates that the target disappears in some frames of the simulated sequence.
@@ III-C Training-test Sets and Annotations
Figure 6: Illustrations of target attributes. (a) shows speed, SCR and swerve times of targets. Middle circle shows target numbers w.r.t. target SCR, and inner and outer circles show the swerving times and target speed of different target SCR levels. Numbers in the pie chart represent the proportion of targets at each SCR level. Numbers in the legends of top-right corner and bottom-right corner represent the proportion of targets in each speed level and swerving times. (b) shows the area and shape distribution of target. “S0-S4” outside the circle represent different target shapes, of which the eccentricity range from $[0,0.2),[0.2,0.4),[0.4,0.6),[0.6,0.8),[0.8,1]$. Lines with different color represent different target area levels. Radius represents the annotation number, and the area under each color line represents the total annotation number of each target area level.
@@ III-D1 Rich Target Diversity
where $\mu_{t}$ is the mean value of image in target region (i.e., yellow area in Fig. 3 (b)). $\mu_{b}$ and $\sigma_{b}$ are the mean and standard deviation value in the local background region (i.e., purple area in Fig. 3 (b)). Statistical values in Fig. 6 (a) show that the number of targets with SCR$<$3 is over a third (i.e., 34.8%), which demonstrates that IRSatVideo-LEO dataset is challenging of not only small (target area$<$$9$$\times$$9$) but also dim targets. In addition, over half of the targets present no swerve and slow speed, while a small number of targets present 1-2 swerving times (46.4%) and medium or fast speed (38.5%), enriching the target diversity and increase the difficulty of IRSatVideo-LEO dataset. In conclusion, our dataset contain high-diversity targets with different numbers, SCR, speed and swerving times, and the distribution of these attributes is controllable and in accordance with the real-world scenario. In addition, Fig. 6 (b) shows the distribution of target area, target shape w.r.t. annotation number. Since target are simulated by Gaussian kernel, we employ eccentricity $e$ for shape distinguish. It can be observed that, target smaller than 10 pixels occupies the largest proportion and most of the targets are in circular shape.
@@ III-D2 Complex Background
where $C_{t}$ represents the background complexity of the $t^{th}$ frame $I_{t}$ in a sequence, and $\bar{C}$ represents the average one of the whole sequence. $s$ represents the gray level in the histogram of $I_{t}$, and $p_{s}(I_{t})$ is the probability of $s$ in $I_{t}$. As shown in Fig. 7 (a), we divide background complexity into four levels: easy $\bar{C}$$\in$$[0,200)$, medium $\bar{C}$$\in$$[200,1000)$, complex $\bar{C}$$\in$$[1000,2000)$ and extreme complex $\bar{C}$$\in$$[2000,\infty)$. It can be observed that our dataset covers a large range of background complexity (from 9 to 5596). In addition, we visualize some example background images in different levels in Fig. 7 (b). It can be observed that easy sequences have low background fluctuation, which increases significantly as the level increases. Moreover, we divide background moving speed (i.e., average shift pixels per frame) into four levels: $\in$$[0,1/10)$, $\in$[1/10,1/3)11013[1/10,1/3)[ 1 / 10 , 1 / 3 ), $\in$$[1/3,1)$ and very $\in$$[1,2]$. It can be observed that there exists all background moving speed levels in each background complexity level, and over 60% sequences are in fast or even higher background moving speed. In conclusion, our dataset contains abundant backgrounds with different complexity and moving speeds.
@@ III-D3 Comparison to Existing IRST Detection Datasets
In this subsection, we make comprehensive comparisons among recent public SIRST datasets [23, 61, 26, 27, 56], MIRST datasets [57, 58, 59, 68, 60] and our IRSatVideo-LEO dataset. As shown in Table II, compared to SIRST datasets, MIRST datasets are in much larger scales (e.g., frame number and target number) and focus on targets with lower SCR (temporal information is necessary for MIRST detection). Note that, many MIRST datasets only provide Bounding Box (BBox) or point annotations, which cannot provide a comprehensive performance evaluation among existing MIRST detection methods (i.e., foreground and background segmentation task). Although IRDST provides mask annotations, the land-based imaging system (e.g., small field of view and disparity variations), and PTZ Camera (e.g., irregular background motion results from fast and flexible platform) cannot meet the actual conditions of satellite videos (e.g., large field of view without disparity variation with respect to remote sensing imaging system, slow & regular background motion with respect to satellite cameras). In conclusion, our dataset is the first large-scale, space-based MIRST dataset with small dim targets & moving background and groundtruth mask annotations, which greatly facilitates the development of IRST detection and tracking in satellite videos.
@@ IV-A Overall Framework
As shown in Fig. 8, the pipeline of our recurrent feature refinement (RFR) framework can be summarized into three components: 1) 3$\times$3 convolution for initial feature extraction. 2) Recurrent feature refinement block (FRB) for recurrent feature alignment, propagation, aggregation and refinement. That is, FRB refine the current feature using recurrently refined previous features, and thus exhibits remarkable efficiency in leveraging long-term temporal dependencies. 3) SIRST detection network for object detection. In this paper, we mainly focus on the design of recurrent FRB, which consists of a pyramid deformable alignment (PDA) module for coarse-to-fine motion compensation, and a temporal-spatial-frequency modulation (TSFM) module for multi-dimensional information refinement. The details of PDA and TSFM are introduced in sections IV-B and IV-C.
@@ IV-C Temporal-Spatial-Frequency Modulation
Temporal attention [52] aims to eliminate the misaligned, occluded, and blur pixels of aligned features for temporal feature aggregation. Specifically, we first calculate the feature similarity between the aligned feature $f_{i-1}^{b}$ and the current feature $g_{i}$ in an embedding space to generate the temporal attention map $M_{i}^{s}$. The process can be defined as:
$M_{i}^{s}=\mathcal{P}_{sigmoid}(\theta(f_{i-1}^{b})^{T}\theta(g_{i})),$ (15)
where $\mathcal{P}_{sigmoid}$ represents sigmoid function that serves as an activator to normalize the temporal attention and stabilize the training process. Awaring of the pixel-level errors, the aligned feature is then multiplied by attention map for temporal information modulation. A plain convolutional layer is also adopted for temporal feature fusion. The process can be defined as:
$f_{i}^{s}=\theta([f_{i-1}^{b}\odot M_{i}^{s},g_{i}]),$ (16)
@@ V-A1 Evaluation Metrics
Area Under Curve: AUC [35] is a the area under ROC, and larger value represents better detection performance.
@@ V-B Comparison to the State-of-the-art Methods
In this subsection, we apply our recurrent feature refinement framework with 5 existing state-of-the-art SIRST detection methods (i.e., ACM [24], ALCNet [25], DNA-Net [26], ISTDU-Net [80] and ResUNet [81]), and make comparisons to the state-of-the-art SIRST detection methods and MIRST detection methods. For SIRST detection methods, we adopt 10 recent state-of-the-art data-driven methods (i.e., ACM [24], ALCNet [25], DNA-Net [26], ISNet [27], UIU-Net [79], RDIAN [60], ISTDU-Net [80], ResUNet [81], RPCANet [33], AGPCNet [32]). For MIRST detection methods, we adopt 7 model-driven methods (i.e., 4DISTD [77], ASTTV_NTLA [20], MSLSTIPT [22], SRSTT [5], IMNN-LWEC[48], NFTDGSTV [54], and RCTVW [78]) and 3 state-of-the-art data-driven methods (i.e., DNANet+DTUM [35], ISTDUNet+DTUM [35] and ResUNet+DTUM [35]). The parameters of 7 model-driven MIRST methods are shown in Table III. For fair comparison, we retrain all the compared data-driven methods on IRSatVideo-LEO dataset from scratch under the same loss function and implemental details. Note that, the training details of SIRST detection methods are inherited from the BasicIRSTD toolbox [85], and the training details of data-driven MIRST detection methods are the same as those in section V-A2. Note that, we adopt a fixed threshold of 0.5 for all the compared algorithms.
@@ V-B1 Quantitative Results
TABLE V: $P_{d}$(×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values achieved by main variants of recurrent framework on MIRST-NUDT-Sea dataset. “-TSWt” represents temporal sliding window framework with temporal length $t$ and stride 1. “-Rt-t” represents recurrent framework with temporal length $t$ for training, and temporal sliding window framework for test with temporal length $t$ and stride $t$. “-Rt-seq” represents recurrent framework with temporal length $t$ for training, and recurrent framework for test with the entire image sequence. “#Params.” represents the number of parameters. FLOPs is computed based on input image sequence with spatial resolution of 256$\times$256 and temporal length of $t$ for “-TSW”, $20$ for “-R”. represents the final model. Best results are shown in boldface.
@@ V-B3 Computational Efficiency
The computational efficiency (the number of parameters, FLOPs, and running time) are evaluated in Table IV. Note that, running time is the total time tested on a 20$\times$1024$\times$1024 image sequence with an Intel(R) Core(TM) i7-8700 CPU @ 3.20GHz, and is averaged over 100 runs. As compared with model-driven methods, data-driven methods achieve large-margin improvements in both detection performance and running time. Among all data-driven methods, MIRST detection methods show superior detection performance with reasonable reasonable increase in computational cost and running time.
TABLE VI: $P_{d}$(×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values achieved by main variants of pyramid deformable alignment on MIRST-NUDT-Sea dataset. “-Flow” represents optical flow-based method for feature alignment. “-Att” represents cross attention module for feature alignment. “-DA” represents deformable convolution for feature alignment. “-PDA” represents pyramid deformable convolution for feature alignment. “#Params.” represents the number of parameters. FLOPs is computed based on input image sequence with a resolution of 20$\times$256$\times$256. ()(*)( ∗ ) represents the final model. Best results are shown in boldface.
@@
Figure 11: Visual results of different alignment methods. (a1), (a2) show the neighborhood and reference images and (b1), (b2) show the corresponding features. (b3) shows the aligned features generated by different alignment methods, and (c2)-(c5) shows the flow (derived by RAFT [86]) between aligned and reference features of ResUNet w/o. (c1) shows the flow between neighborhood and reference features as the baseline results. Flow field color coding scheme is shown in the right. The direction and magnitude of the displacement vector are represented by hue and color intensity, respectively. Flow magnitudes are also presented within each flow.
@@ V-C1 Recurrent Framework
TABLE VII: $P_{d}$(×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values achieved by main variants of temporal-spatial-frequent modulation on MIRST-NUDT-Sea dataset. “-TM” represents temporal attention for feature modulation. “-TSM” represents cascaded temporal attention and spatial attention for feature modulation. “-TSFM” represents cascaded temporal attention, spatial attention, and frequency attention for feature modulation. “#Params.” represents the number of parameters. FLOPs is computed based on input image sequence with a resolution of 20$\times$256$\times$256. represents the final model. Best results are shown in boldface.
@@
  • ResUNet-Rt-t: We employ recurrent framework with temporal length $t$ for training. For test, we employ temporal sliding window of temporal length $t$ and stride $t$ ($t$ image inputs for $t$ image outputs), and recurrent feature refinement is within each temporal sliding window. $t$ is set to 20 in our experiments.
@@
Table V shows the comparative results achieved by recurrent framework and its variants. It can be observed that improvements can be achieved among all variants against the baseline methods, which demonstrate the effectiveness of temporal additional information for IRST detection. All recurrent variants perform superior than temporal sliding window variants for 9.590 $P_{d}$ and 20.402 $F_{a}$ on average, which demonstrates the superiority of recurrent framework. Within the temporal sliding window variants, the detection performance increases first (i.e., 2.055 gain in $P_{a}$ and 5.681 drop in $F_{a}$ from 5 to 7) and then decreases as the number of input frames increases (i.e., 8.316 drop in $P_{a}$ from 7 to 20). This is because, appropriate temporal information can introduce performance improvement, while too much temporal information with inferior fusion methods deteriorates performance instead by ambiguous and misleading information.
Figure 12: Ablation on the temporal modulation (TM) and spatial modulation (SM) of TSFM. We show the temporal attention and spatial attention generated by TM and SM. Features before and after alignment, together with the flow derived by RAFT [86] are also presented. We also present the zoom-in target region for better visualization. Flow field color coding scheme is shown in the right. The direction and magnitude of the displacement vector are represented by hue and color intensity, respectively.
For recurrent variants, due to limited GPU memory and batch-based training process, temporal length of the input video clip is manually pre-defined and fixed for training. During the test phase, the whole sequence (usually contains more than 200 images, and sometimes over 900 images) can be input into the network for recurrent feature refinement through interactions between GPU and CPU. Therefore, we investigate the influence of the input sequence length during training and test phases. It can be observed from Table V that the performance improves steadily as the sequence length increases. This is because, recurrent frameworks effectively use the long-term temporal dependency for substantial performance improvement. Note that, as the number further increases (from 20 to 40), the improvement tends to be saturated. For better trade-off of performance and efficiency, we employ the temporal length of 20 in the final training process. In addition, the $P_{d}$ and $F_{a}$ values tested on recurrent framework with fixed temporal sliding window are inferior to these on recurrent framework with entire sequence (85.726 vs. 86.468 in $P_{d}$ and 19.943 vs. 18.574 in $F_{a}$), which also demonstrate the effectiveness and superiority of recurrent framework.

### v3

@@ III-B2 Target Appearance Simulation
Due to the remote sensing imaging systems and optical diffraction effect, infrared small targets always appear as small points or diffraction spots. As shown in Fig. 3 (a), targets can be approximated by Gaussian kernels $G(h,w,\sigma)$ [78, 79, 68], where $h$, $w$ are major & minor axis, and $\sigma$ represents sigma value. We randomly simulate $N$$\in$$[1,6]$ Gaussian targets for a image sequence, and $h$$\in$$[1,9]$, $w$$\in$$[1,9]$, $\sigma$$\in$$[0.1,1]$. To generate the target appearance sequence of the $n^{th}$ target $G_{nt}(h_{t},w_{t},\sigma_{t})$, as shown in Fig. 1 (b), we randomly set $K$$\in$$[2,5]$ Gaussian kernels $G_{nt_{1}}$, $G_{nt_{2}}$,…,$G_{nt_{K}}$ ($t_{1}$=$1$, $t_{K}$=$T$, $\{t_{2}$$t_{K-1}\}$$\in$$(1,T)$) as the key target appearances in a sequence, and employ quadratic interpolation1 to generate intermediate appearances.
@@ III-B3 Target Trajectory Simulation
Figure 4 shows some examples of target trajectories in satellite video sequences [80]. It can be observed that the real target trajectory is continuous and relatively smooth over a short time interval, which can be approximated using a low-order curve with order $p$$\in$$(0,3)$ [68]. In addition, the target may swerve to change the moving direction and velocity. To simulate the $n^{th}$ target trajectory with $S$$\in$$[0,2]$ swerving times, as shown in Fig. 1 (c), we first synthetic $S$ local target trajectories by low-order curves, and sequentially connect them end-to-end. Then, we employ 1D interpolation to smooth each intersection of two adjacent trajectories for a global smooth trajectory $(x_{nt}^{T},y_{nt}^{T})$. Figure 5 shows the trajectories of two swerved targets in the $1^{st}$ frame reference. It can be observed that target trajectories are locally continuous and globally smooth, which are in accordance with the real-world target motion. It is worth noticing in Fig. 5 (b1) that, the moving direction of one target is opposite to that of the background, which indicates that the target disappears in some frames of the simulated sequence.
@@ III-C Training-test Sets and Annotations
Figure 6: Illustrations of target attributes. (a) shows speed, SCR and swerve times of targets. Middle circle shows target numbers w.r.t. target SCR, and inner and outer circles show the swerving times and target speed of different target SCR levels. Numbers in the pie chart represent the proportion of targets at each SCR level. Numbers in the legends of top-right corner and bottom-right corner represent the proportion of targets in each speed level and swerving times. (b) shows the area and shape distribution of target. “S0-S4” outside the circle represent different target shapes, of which the eccentricity range from $[0,0.2),[0.2,0.4),[0.4,0.6),[0.6,0.8),[0.8,1]$. Lines with different colors represent different target area levels. Radius represents the annotation number, and the area under each color line represents the total annotation number of each target area level.
@@ III-D1 Rich Target Diversity
where $\mu_{t}$ is the mean value of image in target region (i.e., yellow area in Fig. 3 (b)). $\mu_{b}$ and $\sigma_{b}$ are the mean and standard deviation values in the local background region (i.e., purple area in Fig. 3 (b)). Statistical values in Fig. 6 (a) show that the number of targets with SCR$<$3 is over a third (i.e., 34.8%), which demonstrates that IRSatVideo-LEO dataset is challenging of not only small (target area$<$$9$$\times$$9$) but also dim targets. In addition, over half of the targets present no swerve and slow speed, while a small number of targets present 1-2 swerving times (46.4%) and medium or fast speed (38.5%), enriching the target diversity and increase the difficulty of IRSatVideo-LEO dataset. In conclusion, our dataset contains high-diversity targets with different numbers, SCR, speed and swerving times, and the distribution of these attributes is controllable and in accordance with the real-world scenario. In addition, Fig. 6 (b) shows the distribution of target area, target shape w.r.t. annotation number. Since targets are simulated by Gaussian kernel, we employ eccentricity $e$ for shape distinguishing. It can be observed that, targets smaller than 10 pixels occupy the largest proportion, and most of the targets are in circular shape.
@@ III-D2 Complex Background
where $C_{t}$ represents the background complexity of the $t^{th}$ frame $I_{t}$ in a sequence, and $\bar{C}$ represents the average one of the whole sequence. $s$ represents the gray level in the histogram of $I_{t}$, and $p_{s}(I_{t})$ is the probability of $s$ in $I_{t}$. As shown in Fig. 7 (a), we divide background complexity into four levels: easy $\bar{C}$$\in$$[0,200)$, medium $\bar{C}$$\in$$[200,1000)$, complex $\bar{C}$$\in$$[1000,2000)$ and extreme complex $\bar{C}$$\in$$[2000,\infty)$. It can be observed that our dataset covers a large range of background complexity (from 9 to 5596). In addition, we visualize some example background images in different levels in Fig. 7 (b). It can be observed that easy sequences have low background fluctuation, which increases significantly as the level increases. Moreover, we divide background moving speed (i.e., average shift pixels per frame) into four levels: $\in$$[0,1/10)$, $\in$[1/10,1/3)11013[1/10,1/3)[ 1 / 10 , 1 / 3 ), $\in$$[1/3,1)$ and very $\in$$[1,2]$. It can be observed that there exists all background moving speed levels in each background complexity level, and over 60% of sequences are in fast or even higher background moving speed. In conclusion, our dataset contains abundant backgrounds with different complexity and moving speeds.
@@ III-D3 Comparison to Existing IRST Detection Datasets
In this subsection, we make comprehensive comparisons among recent public SIRST datasets [25, 70, 28, 29, 65], MIRST datasets [66, 67, 68, 76, 69] and our IRSatVideo-LEO dataset. As shown in Table II, compared to SIRST datasets, MIRST datasets are in much larger scales (e.g., frame number and target number) and focus on targets with lower SCR (temporal information is necessary for MIRST detection). Note that, many MIRST datasets only provide Bounding Box (BBox) or point annotations, which cannot provide a comprehensive performance evaluation among existing MIRST detection methods (i.e., foreground and background segmentation task). Although IRDST and NUDT-MIRSDT datasets provide mask annotations, the land-based imaging system (e.g., small field of view and disparity variations), and PTZ Camera (e.g., irregular background motion results from fast and flexible platform) or static background motion cannot meet the actual conditions of satellite videos (e.g., large field of view without disparity variation with respect to remote sensing imaging system, slow & regular background motion with respect to satellite cameras). In conclusion, IRSatVideo-LEO is the first large-scale MIRST dataset in satellite videos, that exhibits larger data quantity, higher image resolution, LEO satellite-based moving platform, short wave band imaging system, maneuvering small dim targets, groundtruth mask annotations, making it a valuable resource for research and development in satellite-based video surveillance and remote sensing applications.
@@ IV-A Overall Framework
As shown in Fig. 8, the pipeline of our recurrent feature refinement (RFR) framework can be summarized into three components: 1) 3$\times$3 convolution for initial feature extraction to generate initial feature $g\in\mathbb{R}^{C\times T\times H\times W}$ from the input image sequence $x\in\mathbb{R}^{1\times T\times H\times W}$. 2) Recurrent feature refinement block (FRB) for recurrent feature alignment, propagation, aggregation and refinement. That is, FRB refines the current feature $g_{i}\in\mathbb{R}^{C\times H\times W}$ using recurrently refined previous features $f_{i-1}\in\mathbb{R}^{C\times H\times W}$, and thus exhibits remarkable efficiency in leveraging long-term temporal dependencies. 3) SIRST detection network for object detection. Note that, In this paper, we mainly focus on the design of recurrent FRB, which consists of a pyramid deformable alignment (PDA) module for coarse-to-fine motion compensation, and a temporal-spatial-frequency modulation (TSFM) module for multi-dimensional information refinement. The details of PDA and TSFM are introduced in sections IV-B and IV-C.
@@ IV-C Temporal-Spatial-Frequency Modulation
Temporal attention [54] aims to eliminate the misaligned, occluded, and blur pixels of aligned features for temporal feature aggregation. Specifically, we first calculate the feature similarity between the aligned feature $f_{i-1}^{a}$ and the current feature $g_{i}$ in an embedding space to generate the temporal attention map $M_{i}^{s}$. The process can be defined as:
$M_{i}^{s}=\mathcal{P}_{Sigmoid}(\theta(f_{i-1}^{a})^{T}\theta(g_{i})),$ (15)
where $\mathcal{P}_{Sigmoid}$ represents the Sigmoid function that serves as an activator to normalize the temporal attention and stabilize the training process. Awaring of the pixel-level errors, the aligned feature is then multiplied by attention map for temporal information modulation. A plain convolutional layer is also adopted for temporal feature fusion. The process can be defined as:
$f_{i}^{s}=\theta([f_{i-1}^{a}\odot M_{i}^{s},g_{i}]),$ (16)
@@ V-A1 Evaluation Metrics
Area Under Curve: AUC [37] is the area under ROC, and larger value represents better detection performance.
@@ V-B Comparison to the State-of-the-art Methods
In this subsection, we apply our recurrent feature refinement framework with 5 existing state-of-the-art SIRST detection methods (i.e., ACM [26], ALCNet [27], DNA-Net [28], ISTDU-Net [88] and ResUNet [89]), and make comparisons to the state-of-the-art SIRST detection methods and MIRST detection methods. For SIRST detection methods, we adopt 10 recent state-of-the-art data-driven methods (i.e., ACM [26], ALCNet [27], DNA-Net [28], ISNet [29], UIU-Net [87], RDIAN [69], ISTDU-Net [88], ResUNet [89], RPCANet [35], AGPCNet [34]). For MIRST detection methods, we adopt 7 model-driven methods (i.e., 4DISTD [85], ASTTV_NTLA [22], MSLSTIPT [24], SRSTT [5], IMNN-LWEC[50], NFTDGSTV [63], and RCTVW [86]) and 4 state-of-the-art data-driven methods (i.e., STDMANet [54], DNANet+DTUM [37], ISTDUNet+DTUM [37] and ResUNet+DTUM [37]). The parameters of 7 model-driven MIRST methods are shown in Table III. For fair comparison, we retrain all the compared data-driven methods on IRSatVideo-LEO dataset from scratch under the same loss function and implemental details. Note that, the training details of SIRST detection methods are inherited from the BasicIRSTD toolbox [93], and the training details of data-driven MIRST detection methods are the same as those in section V-A2. Note that, we adopt a fixed threshold of 0.5 for all the compared algorithms.
@@ V-B1 Quantitative Results
TABLE V: $P_{d}$(×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values achieved by main variants of recurrent framework on IRSatVideo-LEO dataset. “-TSWt” represents temporal sliding window framework with temporal length $t$ and stride 1. “-Rt-t” represents recurrent framework with temporal length $t$ for training, and temporal sliding window framework for test with temporal length $t$ and stride $t$. “-Rt-seq” represents recurrent framework with temporal length $t$ for training, and recurrent framework for test with the entire image sequence. “#Params.” represents the number of parameters. FLOPs is computed based on input image sequence with spatial resolution of 256$\times$256 and temporal length of $t$ for “-TSW”, $20$ for “-R”. represents the final model. Best results are shown in boldface.
@@ V-B3 Computational Efficiency
The computational efficiency (the number of parameters, FLOPs, and running time) are evaluated in Table IV. Note that, running time is the total time tested on a 20$\times$1024$\times$1024 image sequence with an Intel(R) Core(TM) i7-8700 CPU @ 3.20GHz, and is averaged over 100 runs. As compared with model-driven methods, data-driven methods achieve large-margin improvements in both detection performance and running time. Among all data-driven methods, MIRST detection methods show superior detection performance with reasonable increase in computational cost and running time.
TABLE VI: $P_{d}$(×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values achieved by main variants of pyramid deformable alignment on IRSatVideo-LEO dataset. “-Flow” represents optical flow-based method for feature alignment. “-Att” represents cross attention module for feature alignment. “-DA” represents deformable convolution for feature alignment. “-PDA” represents pyramid deformable convolution for feature alignment. “#Params.” represents the number of parameters. FLOPs is computed based on input image sequence with a resolution of 20$\times$256$\times$256. ()(*)( ∗ ) represents the final model. Best results are shown in boldface.
@@
Figure 11: Visual results of different alignment methods. (a1), (a2) show the neighborhood and reference images and (b1), (b2) show the corresponding features. (b3) shows the aligned features generated by different alignment methods, and (c2)-(c5) shows the flow (derived by RAFT [94]) between aligned and reference features of ResUNet w/o. (c1) shows the flow between neighborhood and reference features as the baseline results. Flow field color coding scheme is shown on the right. The direction and magnitude of the displacement vector are represented by hue and color intensity, respectively. Flow magnitudes are also presented within each flow.
@@ V-C1 Recurrent Framework
TABLE VII: $P_{d}$(×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values achieved by main variants of temporal-spatial-frequent modulation on IRSatVideo-LEO dataset. “-TM” represents temporal attention for feature modulation. “-TSM” represents cascaded temporal attention and spatial attention for feature modulation. “-TSFM” represents cascaded temporal attention, spatial attention, and frequency attention for feature modulation. “#Params.” represents the number of parameters. FLOPs is computed based on input image sequence with a resolution of 20$\times$256$\times$256. represents the final model. Best results are shown in boldface.
@@
  • ResUNet-Rt-t: We employ recurrent framework with temporal length $t$ for training. For test, we employ a temporal sliding window of temporal length $t$ and stride $t$ ($t$ image inputs for $t$ image outputs), and recurrent feature refinement is within each temporal sliding window. $t$ is set to 20 in our experiments.
@@
Table V shows the comparative results achieved by recurrent framework and its variants. It can be observed that improvements can be achieved among all variants against the baseline methods, which demonstrate the effectiveness of temporal additional information for IRST detection. All recurrent variants perform superior than temporal sliding window variants for 9.590 $P_{d}$ and 20.402 $F_{a}$ on average, which demonstrates the superiority of recurrent framework. Within the temporal sliding window variants, the detection performance increases first (i.e., 2.055 gain in $P_{d}$ and 5.681 drop in $F_{a}$ from 5 to 7) and then decreases as the number of input frames increases (i.e., 8.316 drop in $P_{d}$ from 7 to 20). This is because, appropriate temporal information can introduce performance improvement, while too much temporal information with inferior fusion methods deteriorates performance instead by ambiguous and misleading information.
Figure 12: Ablation on the temporal modulation (TM) and spatial modulation (SM) of TSFM. We show the temporal attention and spatial attention generated by TM and SM. Features before and after alignment, together with the flow derived by RAFT [94] are also presented. We also present the zoom-in target region for better visualization. Flow field color coding scheme is shown on the right. The direction and magnitude of the displacement vector are represented by hue and color intensity, respectively.
For recurrent variants, due to limited GPU memory and batch-based training process, temporal length of the input video clip is manually pre-defined and fixed for training. During the test phase, the whole sequence (usually contains more than 200 images, and sometimes over 900 images) can be input into the network for recurrent feature refinement. Therefore, we investigate the influence of the input sequence length during the training and test phases. It can be observed from Table V that the performance improves steadily as the sequence length increases. This is because, recurrent frameworks effectively use the long-term temporal dependency for substantial performance improvement. Note that, as the number further increases (from 20 to 40), the improvement tends to be saturated. For a better trade-off of performance and efficiency, we employ the temporal length of 20 in the final training process. In addition, the $P_{d}$ and $F_{a}$ values tested on recurrent framework with fixed temporal sliding window are inferior to these on recurrent framework with entire sequence (85.726 vs. 86.468 in $P_{d}$ and 19.943 vs. 18.574 in $F_{a}$), which also demonstrate the effectiveness and superiority of recurrent framework.